40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 25 11:02:39 2009, rana, Summary, PSL, Temperature Box D980400-B.pdf
    Reply  Tue Aug 25 18:42:42 2009, steve, Update, PSL, reference cavity temp box temporarly out of order 
       Reply  Tue Aug 25 21:36:28 2009, Alberto, Update, PSL, reference cavity temp box temporarly out of order 2009-08-25_PSLtrend.png
       Reply  Tue Aug 25 21:55:11 2009, rana, Update, PSL, reference cavity temp box temporarly out of order Untitled.pngrc-heater.jpg
          Reply  Wed Aug 26 15:42:17 2009, Alberto, Update, PSL, reference cavity temp box temporarly out of order 
             Reply  Wed Aug 26 16:11:41 2009, Alberto, Update, PSL, reference cavity temp box temporarly out of order 
                Reply  Wed Aug 26 16:31:34 2009, steve, Update, PSL, reference cavity temp box temporarly out of order 
Message ID: 1943     Entry time: Tue Aug 25 18:42:42 2009     In reply to: 1942     Reply to this: 1945   1946
Author: steve 
Type: Update 
Category: PSL 
Subject: reference cavity temp box temporarly out of order 

Quote:

The PSL Temperature Box (D980400-B-C, what kind of numbering scheme is that?) modified at LHO/LLO ~8 years ago to have better resolution on the in-loop temperature sensors.

I haven't been able to find a DCN / ECN on this, but there's an elog entry from Hugh Radkins here.  I'm also attaching the PDF of the latest drawing (circa 2000) from the DCC.

The schematic doesn't show it, but I am guessing that the T_SENSE inputs are connected to the AD590 chips, and that 4 of these are attached somehow to the RefCav can. IF this is true, I don't understand why there are input resistors on the LT1125 of U1; the AD590 is supposed to be a current source ?

Peter King is supposed to be coming over to work on this today so whoever spots him should force/cajole/entice him to elog what he's done. Film him if necessary.

 

I also think R1-8 should be swapped into metal film resistors for stability. The datasheet says that it puts out 1 uA/K, so the opamps put out 10 mV/K.

J8 and JP1 should be shorted to disable both the tidal and VME control input. Both are unused and a potential source of drift.

 Peter King is updating our temp box as Hugh did at Hanford Oct.22 of 2001 I still have not seen an updated drawing of this.

The LT 1021-7 reference chip will arrive tomorrow morning. This modification should be completed by noon.

 ** The link to the DCN from Hugh is here in the DCC.

ELOG V3.1.3-